OpenStudy (anonymous):

A bug is moving along the parabola y=x^2 at a rate such that its distance from the origin is increasing at a rate of 1 cm/min. At what rates are the x and y coordinates of the bug increasing when it is at the point (2,4)?

OpenStudy (anonymous):

It is a related rates problem. I just don't know how to set up the problem

OpenStudy (dumbcow):

\[y = x^{2}\] \[\frac{dy}{dt} = 2x \frac{dx}{dt}\] letting "d" be distance \[d^{2} =x^{2} +y^{2}\] \[2d \frac{dd}{dt} = 2x \frac{dx}{dt} +2y \frac{dy}{dt}\]

OpenStudy (anonymous):

Then do you just put in the point 2,4 for x and y to solve?

OpenStudy (dumbcow):

yes , find "d" distance to point (2,4) and sub in for dy/dt in distance equation

OpenStudy (anonymous):

so it is squareroot of 6 then?

OpenStudy (dumbcow):

no

OpenStudy (anonymous):

I'm confused then.

OpenStudy (dumbcow):

\[d = \sqrt{2^{2}+4^{2}} = \sqrt{20} = 2\sqrt{5}\] dd/dt is given as 1 \[\rightarrow 2(2\sqrt{5}) = 2(2)\frac{dx}{dt}+2(4)[2(2) \frac{dx}{dt}]\]